I am trying to work out how to go about finding the eact value of
tan(sin^-1(-4/5)

Answer by stanbon(75887)

I am trying to work out how to go about finding the eact value of
tan(sin^-1(-4/5)
-------------
You want the tangent of an angle whose sine is -4/5
By definition of sine, y=-4, r=5
You need x which is sqrt(5^2-(-4)^2) = sqrt(9)=3
Question: is the 3 positive or negative?
The range of sin^-1 is -90 to 90 so the angle we are
talking about must be in the 4th quadrant since y is
negative.
In that quadrant x is positive.
So the tangent you want = y/x = -4/3
---------
Cheers,
Stan H.